# ReferencesSuppReferences

### v4

@@ Abstract
We provide precise conditions for nonparametric identification of causal effects by high-frequency event study regressions, which have been used widely in the recent macroeconomics, financial economics and political economy literatures. The high-frequency event study method regresses changes in an outcome variable on a measure of unexpected changes in a policy variable in a narrow time window around an event or a policy announcement (e.g., a 30-minute window around an FOMC announcement). We show that, contrary to popular belief, the narrow size of the window is not sufficient for identification. Rather, the population regression coefficient identifies a causal estimand when (i) the effect of the policy shock on the outcome does not depend on the other variables (separability) and (ii) the surprise component of the news or event dominates all other variables that are present in the event window (relative exogeneity). Technically, the latter condition requires the ratio between the variance of the policy shock and that of the other variables to be infinite in the event window. Under these conditions, we establish the causal meaning of the event study estimand corresponding to the regression coefficient and the consistency and asymptotic normality of the event study estimator. Notably, this standard linear regression estimator is robust to general forms of nonlinearity. We apply our results to Nakamura and Steinsson’s (2018a) analysis of the real economic effects of monetary policy, providing a simple empirical procedure to analyze the extent to which the standard event study estimator adequately estimates causal effects of interest.
@@ 1 Introduction
When separability and relative exogeneity hold we show that (i) any reverse causality from the outcome variable to the policy variable does not generate bias since the reverse causality is dominated by the policy shock and (ii) common unobserved factors correlated with the policy variable do not generate omitted variables bias since they are also dominated by the policy shock. Since the event study method regresses changes in an outcome on the unexpected changes in a policy variable at dates in the policy sample, the event study estimand is equal to the corresponding population regression coefficient. We establish that the latter identifies a weighted average of standardized marginal causal effects (MCEs) of the policy on the outcome. [^2] The MCE is the derivative of the potential outcome with respect to the policy variable. This should not be confused with the marginal treatment effect (MTE) used in microeconometrics which refers to the treatment effect of a program or intervention for those at the margin of participation [cf. Heckman and Vytlacil (2001)]. When relative exogeneity is violated, we show that the event study estimand can be decomposed into the same weighted average of MCEs and a selection bias factor. The selection bias factor is decreasing in the variance of the policy shock so that even when relative exogeneity fails, the event study estimator will not have substantial bias so long as the variance of the policy shock is relatively large.
@@
It is difficult to use statistical tests to verify relative exogeneity. Existing tests for (in)finite variance [e.g., Trapani (2016)] require a large sample of the corresponding random variable to be observed or residuals from a correctly-specified regression. Given the potential for simultaneity and omitted variables bias in high-frequency event study regressions, these requirements are not satisfied in the current context. In addition, the OLS event study estimator can still perform well when relative exogeneity holds approximately. A test for infinite variance of the policy shocks alone, even when feasible, is unable to detect when relative exogeneity provides a good enough approximation for the event study estimator to perform well in practice. Instead of trying to test for infinite variance, we introduce a simple empirical procedure that can be used for a sensitivity analysis to diagnose whether relative exogeneity is “close enough” to holding that the OLS event study estimator should be expected to have mean-squared error at least as small as that of an oracle estimator in a corresponding regression with no endogeneity. This sensitivity analysis uses information available in the relevant control sample, e.g., days when there is no FOMC meeting, in order to obtain a proxy for the variance of the variables that are not specific to the policy within the event window. This procedure requires no additional data relative to that already used by existing high-frequency event studies and can thus be applied easily. We introduce our procedure in the context of an empirical example aimed at assessing the causal effects of monetary policy news on real interest rates. The empirical results show that relative exogeneity is likely to approximately hold in the analysis of Nakamura and Steinsson (2018a) based on a 30-minute or 1-day window for the outcome and policy variable.
@@
Finally, we use our identification framework to shed light on the recent debate in the literature on some puzzling event study regression results involving the causal effect of monetary policy on Blue Chip forecasts of real GDP. We show that the estimates with signs opposite to the predictions from standard macroeconomic models likely arise because relative exogeneity fails in this type of regression. Since the Blue Chip forecast revision is constructed as a one-month change while the monetary policy surprise is constructed as a 30-minute change, it is highly unlikely that the variance of the policy shocks dominates the variance of the other variables that determine the Blue Chip forecast: the latter aggregate news over a much longer time frame. Thus, any endogeneity of the policy surprise is not likely to be drowned out by the variation in the policy shock.
@@
The rest of the paper is structured as follows. In Section 2 we conduct a review of the high-frequency event study literature and detail a few empirical examples, present our identification results and discuss robustness to information leakage. Section 3 presents the asymptotic properties of the OLS event study estimator. Section 4 proposes a sensitivity analysis to verify the identification conditions in the context of an empirical example and discusses some identification issues in event study regressions that involve Blue Chip forecasts. We conclude in Section 5.
@@ 2.1 Event Study Design
[Macroeconomic announcements: Faust, Rogers, Wang, and Wright (2007), Gürkaynaky, Kısacıkoçlu, and Wright (2020) and Gürkaynak, Sack, and Swanson (2005)] Several works studied the effects of macroeconomic announcements on changes in asset prices and exchange rates in a narrow window around news releases such as the U.S. employment report, U.S. GDP releases, Census reports on retail sales, and CPI and PPI data releases. Gürkaynak, Sack, and Swanson (2005) also estimated the event study regression (2.2) but with a vector-valued $D_{t}$ and coefficient $\beta$, where $Y_{t}$ is the log-return on an asset or a change in a bond yield and $D_{t}$ is a vector of news, or unexpected, components of the considered macroeconomic announcements and $u_{t}$ is a serially uncorrelated error term. [^5] Actually, Gürkaynak, Sack, and Swanson (2005) considered a vector of asset prices $Y_{t}$. But since the cross-sectional variation is not exploited for identification we simply take $Y_{t}$ to be a scalar. Unlike for the FOMC announcements, there are no traded instruments from which to infer market expectations for macroeconomic announcements. Thus, in order to identify surprise announcements the literature relies instead on economists’ forecasts from surveys: each element of $D_{t}$ is computed as the difference between the actual macroeconomic data release and its market expectation obtained from the most recent survey. These surveys are the Blue Chip Economic Indicators Survey or those run by Action Economics, or alternatively by Bloomberg. Thus, $D_{t}$ captures the surprise component of each data release.
@@ 2.2 Nonparametric Event Study Design and Identification Conditions
and refer to this limiting case as relative exogeneity. [^10] More precisely, relative exogeneity holds when the ratios between the variance of the policy shock and the variances of the other variables in the event window diverge. Thus, one may instead frame relative exogeneity as the condition for which the variance of the policy shock is finite and the variance of the background noise vanishes in the event window. The same identification results can also be shown to hold under this latter condition. The condition (2.4) makes clear one point that has been overlooked by the empirical literature. Namely, it requires that the “large” variance condition for the policy shock has to hold for all $t\in\mathbf{P}$. For example, suppose that it is satisfied only for a few announcements in the policy sample. Since the variance cannot be negative, an estimate of the average variance computed in the policy sample could still be very large. This could be misleadingly interpreted as support for relative exogeneity. However, the endogeneity in most of the policy sample would not be overcome and relative exogeneity would fail. Hence, verifying that the sample variance of $D_{t}$ is large does not allow one to conclude that relative exogeneity holds.
@@ 2.3 Identification Results
Note that the high-frequency event study method is different from heteroskedasticity-based identification [cf. rigobon:2003], sometimes referred to as Rigobon’s method. The latter also uses information from the control sample, which includes windows (either 30-minute or 1-day windows) that do not bracket an announcement or event. The main heteroskedasticity-based identification condition is that the volatility of the policy shock is larger in the event windows than in the control windows. This is different from relative exogeneity for two reasons. First, the required increase in volatility of the policy shock is not relative to the other variables in the event window but relative to the policy shock in the control windows. Second, the increase in volatility need not be infinite. Thus, heteroskedasticity-based identification requires neither stronger nor weaker conditions than the relative exogeneity condition. In addition, Rigobon’s estimator is an instrumental variables estimator and therefore different from the OLS high-frequency event study estimator. Since these two estimators correspond to different estimands, they identify different causal effect in general.
@@ 2.4 Relation of the Theoretical Results with the Empirical Literature
There are essentially no formal identification results about high-frequency event studies in the literature. Empirical works have often mentioned that changes in the policy variable in the event windows are dominated by the information about future monetary policy contained in the FOMC announcements [see, e.g., Nakamura and Steinsson (2018a)]. However, precise conditions have not been provided. Rigobon and Sack (2004) noted that the bias of the OLS event study estimator disappears in a stylized linear model when relative exogeneity holds. However, the model they analyzed is admittedly a “clear oversimplification”. As such, this stylized model does not enable practically-relevant identification analysis or econometric results. In addition, showing that the bias disappears in this context does not imply identification of a causal estimand. Nevertheless, we credit Rigobon and Sack’s heuristic analysis as an early insight into understanding the validity of the OLS event study estimator. Gürkaynak and Wright (2013) surveyed the empirical literature and presented useful discussions. They also argued that the event study regression works even when the variance of the policy shock is not large relative to the variance of the other variables because in a narrow window around an FOMC announcement the policy news can depend on lagged changes in asset returns but not on contemporaneous changes. That is, they wrote $D_{t}=\alpha Y_{t-j}+Z^{\prime}_{t-j}\gamma_{2}+e_{t}$, where $j\geq 1$. However, as we explain in Example 1, the choice of a narrow window is not sufficient for ruling out omitted variables bias. Further, when $Y_{t}$ is the change in an asset price and $D_{t}$ is, for example, the price change in the Federal funds or eurodollar futures, assuming that $Y_{t}$ does not contemporaneously affect $D_{t}$ constitutes a strong empirical restriction.
@@ 3 Properties of the OLS Event Study Estimator
$J$ $=\int_{0}^{1}J\left(s\right)\mathrm{d}s,\qquad\mathrm{and}\qquad J\left(s\right)=\lim_{T_{P}\rightarrow\infty}\mathbb{E}\left[\varepsilon_{\left\lfloor T_{P}s\right\rfloor}^{2}\left(D_{\left\lfloor T_{P}s\right\rfloor}^{*}-\mathbb{E}(D_{\left\lfloor T_{P}s\right\rfloor}^{*}\right)^{2}\right].$
@@
$J$ $=\lim_{T_{P}\rightarrow\infty}\frac{1}{T_{P}}\sum_{t=1}^{T_{P}}\sum_{l=1}^{T_{P}}\mathbb{E}\left[\varepsilon_{t}\varepsilon_{l}\left(D_{t}^{*}-\mathbb{E}(D_{t}^{*}\right)\left(D_{l}^{*}-\mathbb{E}(D_{l}^{*}\right)\right]$
$=\int_{0}^{1}\int_{0}^{1}\lim_{T_{P}\rightarrow\infty}\mathbb{E}\left[\varepsilon_{\left\lfloor T_{P}s\right\rfloor}\varepsilon_{\left\lfloor T_{P}r\right\rfloor}\left(D_{\left\lfloor T_{P}s\right\rfloor}^{*}-\mathbb{E}(D_{\left\lfloor T_{P}s\right\rfloor}^{*}\right)\left(D_{\left\lfloor T_{P}r\right\rfloor}^{*}-\mathbb{E}(D_{\left\lfloor T_{P}r\right\rfloor}^{*}\right)\right]\mathrm{d}s\mathrm{d}r.$
@@
Theorem 5 shows that heteroskedasticity-robust standard errors suffice for valid inference when there is no serial correlation in $u_{t}$, $e_{t}$ and $Z_{t}$. Under covariance-stationarity $J\left(s\right)$ does not depend on $s,$ and so $J=\mathbb{E}[\varepsilon_{t}^{2}\left(D_{t}^{*}-\mathbb{E}(D_{t}^{*}\right)^{2}]$. However, heteroskedasticity-robust standard errors are still required as $\varepsilon_{t}$ may not be homoskedastic. Finally, the following corollary establishes the unbiased asymptotic normality of $\widehat{\beta}_{\mathrm{ES}}$ under relative exogeneity.
@@
Let Assumptions 1-6 and 8-9 hold. Then as $T_{P}\rightarrow\infty$, then $\min_{t\in\mathbf{P}}\sigma_{e,t}^{2}\rightarrow\infty$,

### v5

@@ Abstract
We consider identification, estimation and inference in high-frequency event study regressions, which have been used widely in the recent macroeconomics, financial economics and political economy literatures. The high-frequency event study method regresses changes in an outcome variable on a measure of unexpected changes in a policy variable in a narrow time window around an event or a policy announcement (e.g., a 30-minute window around an FOMC announcement). We show that, contrary to popular belief, the narrow size of the window is not sufficient for identification. Rather, the population regression coefficient identifies a causal estimand when (i) the effect of the policy shock on the outcome does not depend on the other variables (separability) and (ii) the surprise component of the news or event dominates all other variables that are present in the event window (relative exogeneity). Technically, the latter condition requires the ratio between the variance of the policy shock and that of the other variables to be infinite in the event window. Under these conditions, we establish the causal meaning of the event study estimand corresponding to the regression coefficient and super-consistency of the event study estimator with rate of convergence faster than the parametric rate. We show the asymptotic normality of the estimator and propose bias-corrected inference. We also provide bounds on the worst-case bias and use them to quantify its impact on the worst-case coverage properties of confidence intervals, as well as to construct a bias-aware critical value. Notably, this standard linear regression estimator is robust to general forms of nonlinearity. We apply our results to Nakamura and Steinsson’s (2018a) analysis of the real economic effects of monetary policy, providing a simple empirical procedure to analyze the extent to which the standard event study estimator adequately estimates causal effects of interest.
@@ 1 Introduction
When separability and relative exogeneity hold we show that (i) any reverse causality from the outcome variable to the policy variable does not generate bias since the reverse causality is dominated by the policy shock and (ii) common unobserved factors correlated with the policy variable do not generate omitted variables bias since they are also dominated by the policy shock. Since the event study method regresses changes in an outcome on the unexpected changes in a policy variable at dates in the policy sample, the event study estimand is equal to the corresponding population regression coefficient. We establish that the latter identifies a weighted average of standardized marginal causal effects (MCEs) of the policy on the outcome. [^2] The MCE is the derivative of the potential outcome with respect to the policy variable. This should not be confused with the marginal treatment effect (MTE) used in microeconometrics which refers to the treatment effect of a program or intervention for those at the margin of participation [cf. Heckman and Vytlacil (2001)]. When relative exogeneity is violated, we show that the event study estimand can be decomposed into the same weighted average of MCEs and a selection bias factor. The magnitude of the selection bias factor is decreasing in the variance of the policy shock so that even when relative exogeneity fails, the event study estimator will not have substantial bias so long as the variance of the policy shock is relatively large.
@@
It is difficult to use statistical tests to verify relative exogeneity. Existing tests for (in)finite variance [e.g., Trapani (2016)] require a large sample of the corresponding random variable to be observed or residuals from a correctly-specified regression. Given the potential for simultaneity and omitted variables bias in high-frequency event study regressions, these requirements are not satisfied in the current context. In addition, the OLS event study estimator can still perform well when relative exogeneity holds approximately. A test for infinite variance of the policy shocks alone, even when feasible, is unable to detect when relative exogeneity provides a good enough approximation for the event study estimator to perform well in practice. Instead of trying to test for infinite variance, we introduce a simple empirical procedure that can be used for a sensitivity analysis to diagnose whether relative exogeneity is “close enough” to holding that the OLS event study estimator should be expected to have mean-squared error at least as small as that of an oracle estimator in a corresponding regression with no endogeneity. This sensitivity analysis uses information available in the relevant control sample, e.g., days when there is no FOMC meeting, in order to obtain a proxy for the variance of the variables that are not specific to the policy within the event window. The procedure can accommodate nonlinearities, heteroskedasticity and serial correlation. This procedure requires no additional data relative to that already used by existing high-frequency event studies and can thus be applied easily. We introduce our procedure in the context of an empirical example aimed at assessing the causal effects of monetary policy news on real interest rates. The empirical results show that relative exogeneity is likely to approximately hold in the analysis of Nakamura and Steinsson (2018a) based on a 30-minute or 1-day window for the outcome and policy variable.
Under separability and relative exogeneity, we establish the super-consistency of the OLS event study estimator for a weighted average of standardized MCEs, with a rate of convergence equal to the square root of the sample size multiplied by the standard deviation of the policy variable. Asymptotic bias arises when relative exogeneity is not “strong enough”—that is, when the standard deviation of the policy variable diverges at the same rate or more slowly than the square root of the sample size. Nonetheless, we show that under the “sharp null” hypothesis of zero MCEs, it is possible to consistently estimate the asymptotic bias and develop inference refinements based on bias correction, even when relative exogeneity is not “strong enough”.
Since the bias vanishes only in the limit under relative exogeneity, we consider methods to quantify how this bias can affect the worst-case properties of the estimator by deriving a bound on the worst-case bias. In particular, we use this bound to study the worst-case asymptotic coverage of standard confidence intervals based on the OLS event study estimator. We then propose a bias-aware critical value that accounts for the estimator’s potential asymptotic bias in addition to its variance, by adjusting the standard normal critical value upward to compensate for the bias. The resulting bias-aware inference does not rely on the imposition of the “sharp null” hypothesis.
@@
We use our identification framework to shed light on the recent debate in the literature on some puzzling event study regression results involving the causal effect of monetary policy on Blue Chip forecasts of real GDP. We show that the estimates with signs opposite to the predictions from standard macroeconomic models likely arise because relative exogeneity fails in this type of regression. Since the Blue Chip forecast revision is constructed as a one-month change while the monetary policy surprise is constructed as a 30-minute change, it is highly unlikely that the variance of the policy shocks dominates the variance of the other variables that determine the Blue Chip forecast: the latter aggregate news over a much longer time frame. Thus, any endogeneity of the policy surprise is not likely to be drowned out by the variation in the policy shock.
Before proceeding with the analysis, it is worth clarifying the estimation approach we analyze in relation to other related approaches. We analyze the widely popular event study estimand in order to determine the causal effects under study in empirical practice when estimating high-frequency event study regressions. Under the conditions we provide, the high-frequency event study estimand identifies a weighted average of MCEs of the policy variable rather than those of exogenous shocks, which are more customary in impulse response analysis of structural vector autoregressions (SVARs). Our approach is thus to elucidate which causal effects are being analyzed in high-frequency event study regressions, not to take a stand on which type of MCEs are more or less policy-relevant. However, we note that there are many applications for which identifying the causal effects of the policy variable is of interest. These include, for example, estimating the slope of demand functions, price elasticities, and response or reaction coefficients—such as how asset prices respond to monetary policy. High-frequency policy surprises used in event studies are also employed as instruments for the policy shock in SVAR contexts to identify impulse responses. This approach corresponds to a different estimand and identified causal effect, which are not the focus of the present paper.
Although it bares some conceptual similarity, the high-frequency event study approach is distinct from regression discontinuity. The former involves a single observation for each announcement or event, and the estimand is the linear projection coefficient, given by the ratio of the covariance between the outcome and the policy variable to the variance of the policy variable. In contrast, regression discontinuity relies on many observations to the left and right of a given cutoff, and the estimand is the difference in the conditional mean of the outcome at the cutoff. A key identification condition for regression discontinuity is local continuity of the mean of the potential outcome in the absence of treatment around the cutoff. The high-frequency event study estimator does not converge to the regression discontinuity estimand, and local continuity is therefore not sufficient to recover a causal effect—even if one assumes the announcement window shrinks to zero. Nevertheless, the infinite variance condition required by relative exogeneity imposes a discontinuity in the policy variable around the announcement, which is similar in spirit to the discontinuity used in regression discontinuity designs. The technical difference is that, in high-frequency identification, the bias is scaled by the inverse of the ratio of the variance of the policy shock to the variance of other shocks, which converges to zero. In contrast, in regression discontinuity, the bias is eliminated through local differencing.
@@
The rest of the paper is structured as follows. Section 2 reviews the high-frequency event study literature, presents several empirical examples, introduces our identification results, and discusses robustness to information leakage. Section 3 establishes the asymptotic properties of the OLS event study estimator. Section 4 derives bounds on the asymptotic bias, examines the worst-case asymptotic coverage properties of confidence intervals based on the OLS event study estimator, and proposes bias-aware inference. Section 5 develops a sensitivity analysis to assess the identification conditions, applies both the sensitivity analysis and our inference results to an empirical example, and discusses identification issues in event study regressions involving Blue Chip forecasts. Section 6 concludes.
@@ 2.1 Event Study Design
[Macroeconomic announcements: Faust, Rogers, Wang, and Wright (2007), Gürkaynaky, Kısacıkoçlu, and Wright (2020), Gürkaynak, Sack, and Swanson (2005) and kanzig:2021 (kanzig:2021; kanzig:2021_Carbon)] Several works studied the effects of macroeconomic announcements on changes in asset prices and exchange rates in a narrow window around news releases such as the U.S. employment report, U.S. GDP releases, Census reports on retail sales, and CPI and PPI data releases. Gürkaynak, Sack, and Swanson (2005) also estimated the event study regression (2.2) but with a vector-valued $D_{t}$ and coefficient $\beta$, where $Y_{t}$ is the log-return on an asset or a change in a bond yield and $D_{t}$ is a vector of news, or unexpected, components of the considered macroeconomic announcements and $u_{t}$ is a serially uncorrelated error term. [^5] Actually, Gürkaynak, Sack, and Swanson (2005) considered a vector of asset prices $Y_{t}$. But since the cross-sectional variation is not exploited for identification we simply take $Y_{t}$ to be a scalar. Unlike for the FOMC announcements, there are no traded instruments from which to infer market expectations for macroeconomic announcements. Thus, in order to identify surprise announcements the literature relies instead on economists’ forecasts from surveys: each element of $D_{t}$ is computed as the difference between the actual macroeconomic data release and its market expectation obtained from the most recent survey. These surveys are the Blue Chip Economic Indicators Survey or those run by Action Economics, or alternatively by Bloomberg. Thus, $D_{t}$ captures the surprise component of each data release.
@@ 2.2 Nonparametric Event Study Design and Identification Conditions
and refer to this limiting case as relative exogeneity. [^10] More precisely, relative exogeneity holds when the ratios between the variance of the policy shock and the variances of the other variables in the event window diverge. Thus, one may instead frame relative exogeneity as the condition for which the variance of the policy shock is finite and the variance of the background noise vanishes in the event window in analogy with a continuous time jump-diffusion model being sampled at high frequency, where the diffusion component corresponds to background noise and the jump component corresponds to the policy shock. See, e.g., bandi/nguyen:2003. We show that the same identification results hold under this latter condition in Appendix B. The condition (2.4) makes clear one point that has been overlooked by the empirical literature. Namely, it requires that the “large” variance condition for the policy shock has to hold for all $t\in\mathbf{P}$. For example, suppose that it is satisfied only for a few announcements in the policy sample. Since the variance cannot be negative, an estimate of the average variance computed in the policy sample could still be very large. This could be misleadingly interpreted as support for relative exogeneity. However, the endogeneity in most of the policy sample would not be overcome and relative exogeneity would fail. Hence, verifying that the sample variance of $D_{t}$ is large does not allow one to conclude that relative exogeneity holds.
@@ 2.3 Identification Results
Note that the high-frequency event study method is different from heteroskedasticity-based identification [cf. lewis:2020 and Rigobon (2003)], sometimes referred to as Rigobon’s method. The latter also uses information from the control sample, which includes windows (either 30-minute or 1-day windows) that do not bracket an announcement or event. The main heteroskedasticity-based identification condition is that the volatility of the policy shock is larger in the event windows than in the control windows. This is different from relative exogeneity for two reasons. First, the required increase in volatility of the policy shock is not relative to the other variables in the event window but relative to the policy shock in the control windows. Second, the increase in volatility need not be infinite. Thus, heteroskedasticity-based identification requires neither stronger nor weaker conditions than the relative exogeneity condition. In addition, Rigobon’s estimator is an instrumental variables estimator and therefore different from the OLS high-frequency event study estimator. Since these two estimators correspond to different estimands, they identify different causal effect in general.
@@ 2.4 Relation of the Theoretical Results with the Empirical Literature
There are essentially no formal identification results about high-frequency event studies in the literature. Empirical works have often mentioned that changes in the policy variable in the event windows are dominated by the information about future monetary policy contained in the FOMC announcements [see, e.g., Nakamura and Steinsson (2018a)]. However, precise conditions have not been provided. Rigobon and Sack (2004) noted that the bias of the OLS event study estimator disappears in a stylized linear model when relative exogeneity holds. However, this result does not suffice to prove identification of a causal estimand. Further, the model they analyzed is admittedly a “clear oversimplification”. As such, this stylized model does not enable practically-relevant identification analysis or econometric results. Nevertheless, we credit Rigobon and Sack’s heuristic analysis as an early insight into understanding the validity of the OLS event study estimator. Gürkaynak and Wright (2013) surveyed the empirical literature and presented useful discussions. They also argued that the event study regression works even when the variance of the policy shock is not large relative to the variance of the other variables because in a narrow window around an FOMC announcement the policy news can depend on lagged changes in asset returns but not on contemporaneous changes. That is, they wrote $D_{t}=\alpha Y_{t-j}+Z^{\prime}_{t-j}\gamma_{2}+e_{t}$, where $j\geq 1$. However, as we explain in Example 1, the choice of a narrow window is not sufficient for ruling out omitted variables bias. Further, when $Y_{t}$ is the change in an asset price and $D_{t}$ is, for example, the price change in the Federal funds or eurodollar futures, assuming that $Y_{t}$ does not contemporaneously affect $D_{t}$ constitutes a strong empirical restriction.
@@ 3 Properties of the OLS Event Study Estimator
$J$ $=\int_{0}^{1}J\left(s\right)\mathrm{d}s,\qquad\mathrm{and}\qquad J\left(s\right)=\lim_{T_{P}\rightarrow\infty}\mathbb{E}\left[\varepsilon_{\left\lfloor T_{P}s\right\rfloor}^{2}\left(D_{\left\lfloor T_{P}s\right\rfloor}^{*}-\mathbb{E}(D_{\left\lfloor T_{P}s\right\rfloor}^{*})\right)^{2}\right].$
@@
$J$ $=\lim_{T_{P}\rightarrow\infty}\frac{1}{T_{P}}\sum_{t=1}^{T_{P}}\sum_{l=1}^{T_{P}}\mathbb{E}\left[\varepsilon_{t}\varepsilon_{l}\left(D_{t}^{*}-\mathbb{E}(D_{t}^{*})\right)\left(D_{l}^{*}-\mathbb{E}(D_{l}^{*})\right)\right]$
$=\int_{0}^{1}\int_{0}^{1}\lim_{T_{P}\rightarrow\infty}\mathbb{E}\left[\varepsilon_{\left\lfloor T_{P}s\right\rfloor}\varepsilon_{\left\lfloor T_{P}r\right\rfloor}\left(D_{\left\lfloor T_{P}s\right\rfloor}^{*}-\mathbb{E}(D_{\left\lfloor T_{P}s\right\rfloor}^{*})\right)\left(D_{\left\lfloor T_{P}r\right\rfloor}^{*}-\mathbb{E}(D_{\left\lfloor T_{P}r\right\rfloor}^{*})\right)\right]\mathrm{d}s\mathrm{d}r.$
@@
Theorem 5 shows that heteroskedasticity-robust standard errors suffice for valid inference when there is no serial correlation in $u_{t}$, $e_{t}$ and $Z_{t}$. Under covariance-stationarity $J\left(s\right)$ does not depend on $s,$ and so $J=\mathbb{E}[\varepsilon_{t}^{2}\left(D_{t}^{*}-\mathbb{E}(D_{t}^{*})\right)^{2}]$. However, heteroskedasticity-robust standard errors are still required as $\varepsilon_{t}$ may not be homoskedastic.
Theorem 5 establishes the asymptotic normality of $\hat{\beta}_{\mathrm{ES}}$ as an estimator for the biased quantity $\beta_{\mathrm{ES}}$. To conduct unbiased inference on the weighted average of standardized MCEs in (3.1) using $\widehat{\beta}_{\mathrm{ES}}$, we require a strengthening of relative exogeneity as formalized in the following corollary.
@@
Let Assumptions 1-6 and 8-9 hold. Then if $\sqrt{T_{P}}/\sigma_{D}\rightarrow c\in[0,\infty)$ as $T_{P}\rightarrow\infty$,